OpenStudy (anonymous):

Nevaeh finances $147,500 to purchase a townhouse with a 30/8 balloon mortgage at 3.6%. What is the remaining balance at the end of the mortgage? $83,122.40 $93,916.00 $113,753.73 $122,166.82

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@wio

OpenStudy (plasmataco):

hmm...

OpenStudy (plasmataco):

what do you mean by 30/8 baloon morgage?

OpenStudy (anonymous):

The 30 means that the beginning fixed monthly payments are going to be calculated as if the mortgage would be paid off in 30 years. The 8 means that fixed payments will be made for 7 years

OpenStudy (anonymous):

sorry for 8 years

OpenStudy (plasmataco):

147000*(3.6)^30?

OpenStudy (anonymous):

@zepdrix

OpenStudy (plasmataco):

there is a formula for this.

OpenStudy (anonymous):

@mathmate can you help me

OpenStudy (anonymous):

hello @mathmate

OpenStudy (anonymous):

are you there @mathmate

OpenStudy (anonymous):

@hero can you help me out here

OpenStudy (mathmate):

Yeah, I am on something. I will help you if no one does, a little later on!

OpenStudy (anonymous):

oh ok @mathmate

OpenStudy (anonymous):

@jim_thompson5910 right?

jimthompson5910 (jim_thompson5910):

let me think

jimthompson5910 (jim_thompson5910):

ok a calculator is best for this go here http://www.bankrate.com/calculators/mortgages/mortgage-calculator.aspx and type in what you see in the attached image I'm posting

jimthompson5910 (jim_thompson5910):

instead of hitting "calculate" hit the "show amortization schedule" button which is below the calculate button let me know when you've done that

OpenStudy (anonymous):

alright did it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok so 8 years from now is 2015+8 = 2023 scroll down til you get to the row that starts with June 2023

jimthompson5910 (jim_thompson5910):

what is the dollar figure in the last column of that row?

OpenStudy (anonymous):

$122,166.82

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Aaliyah obtained a 25-year, fixed rate mortgage for $220,500 on a home that cost her $243,950. If the interest rate on the mortgage is 6.15%, how much interest will Aaliya pay over the life of the mortgage?

OpenStudy (anonymous):

Would this one be the same thing?

jimthompson5910 (jim_thompson5910):

that's the remaining balance after 8 years that amount is to be paid off or refinanced somehow

jimthompson5910 (jim_thompson5910):

you can use the same calculator, yes

OpenStudy (anonymous):

wait but it gives me two values @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

because it says that she is recieving a fixed rate

OpenStudy (anonymous):

that she got at first for 243950

OpenStudy (anonymous):

do i plug in for the mortgage value 220,500

jimthompson5910 (jim_thompson5910):

hmm not sure, but let me think

OpenStudy (anonymous):

yeah thats what got me confused at the beginning

jimthompson5910 (jim_thompson5910):

does it give you any answer choices?

OpenStudy (anonymous):

$257,766 $188,341 $234,316 $211,791

jimthompson5910 (jim_thompson5910):

ok you're going to type in 220500 into the "Mortgage amount" box

jimthompson5910 (jim_thompson5910):

6.15 is the interest rate 25 is the number of years use that calculator to figure out the monthly payment and tell me what you get

OpenStudy (anonymous):

I got 1,440.97

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's the monthly payment multiply that by 300 300 months = 25 years * (12 mon/1 year)

OpenStudy (anonymous):

so that would be 432291

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

now subtract off the mortgaged amount 432,291-220,500 = 211,791

OpenStudy (anonymous):

Riley made a 18% down payment on a home that cost $212,400. He financed the remaining amount using a 25-year, fixed-rate mortgage at 5.85%. His monthly payment will be $1,106.25. Riley will pay for two discount points, a 0.5% origination fee, brokerage fee, state documentary taxes on the deed and the mortgage, and the intangible tax. Discount points equal 1% of the mortgage amount. Documentary stamp tax on deed is $0.70 per $100 or portion thereof. Documentary stamp tax on mortgage is $0.35 per $100 or portion thereof. Mortgage broker fee is $185 plus 3% of the mortgage amount. Intangible tax is 0.2% of the mortgage amount. What is the total cost of Riley’s home including the principal, interest, down payment, and fees? $382,351.08 $556,484.08 $384,629.00 $346,767.20

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let me think, but mathmate go ahead and post what you get

OpenStudy (mathmate):

@gabylovesu Sorry, it's getting late for me tonight. If the calculator does not do your problem, you can tag me or @jim_thompson5910 tomorrow. But I think if the calculator worked for the first problem, it should be good for similar problems.

jimthompson5910 (jim_thompson5910):

still thinking but I may have the answer, one sec

OpenStudy (anonymous):

???

jimthompson5910 (jim_thompson5910):

I'm stuck on the stamp parts. Do you have any formulas for those? I'm trying to determine if the stamp taxes are based on the mortgage amount or the value of the home.

OpenStudy (anonymous):

Like the documentary stamp act

OpenStudy (anonymous):

???

jimthompson5910 (jim_thompson5910):

yeah for the deed and the mortgage

jimthompson5910 (jim_thompson5910):

do you see those formulas anywhere in your notes?

OpenStudy (anonymous):

well they are not really formulas but steps

jimthompson5910 (jim_thompson5910):

can I have a look? through a screenshot maybe?

OpenStudy (anonymous):

This is an example of how they used it The buyer typically pays the documentary stamp tax on the mortgage. In Florida, a buyer will pay $0.35 for every $100 (or portion thereof) of borrowed money. To determine the tax on the mortgage, round the amount borrowed up to the next $100. Then multiply that amount by 0.35 divided by 100, or 0.0035.

jimthompson5910 (jim_thompson5910):

ok I'll try that

jimthompson5910 (jim_thompson5910):

I'm off by about $35 and I'm not sure why

jimthompson5910 (jim_thompson5910):

this is what I'm getting Downpayment = 0.18*212400 = 38232 Mortgage Amount = 212400 - 38232 = 174168 Total amount paid back (principal + interest) = 1106.25*25*12 = 331875 Discount points = 2*0.01*174168 = 3483.36 Origination Fee = 0.005*174168 = 870.84 Documentary stamp tax on deed (based on home value) 212,400*0.70/100 = 1486.80 Documentary stamp tax on mortgage (based on mortgage value) 174168 rounds to 174200 174200*0.35/100 = 609.7 Mortgage broker fee = 185 + 0.03*174168 = 5410.04 Intangible Tax = 0.002*174168 = 348.336 = 348.34 Add up the results: 38232+331875+3483.36+870.84+1486.80+609.7+5410.04+348.34 = 382,316.08

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

is there anything in your notes about "Documentary stamp tax on deed" ?

jimthompson5910 (jim_thompson5910):

you provided the "documentary stamp tax on the mortgage" portion and that was helpful

OpenStudy (anonymous):

Yeah I do @jim_thompson5910

OpenStudy (anonymous):

The seller typically pays the documentary stamp tax on the deed. As one example, in almost every county in Florida, the seller is charged $0.70 for every $100 (or portion thereof) on the sale price of the house. To determine the tax on the deed, round the selling price up to the next $100. Then multiply that amount by 0.70 divided by 100, or 0.007. Regardless of whether a buyer pays cash for a house or uses financing, the seller is still responsible for the documentary stamp tax on the selling price.

jimthompson5910 (jim_thompson5910):

ok let me try something

jimthompson5910 (jim_thompson5910):

yeah I'm still getting 382,316.08 so I'm guessing the answer should be 382,351.08 there's either a rounding error somewhere or I'm missing some fee(s) ?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

honestly i don't know why they make it so complicated with all these taxes and fees. It's kinda ridiculous

OpenStudy (anonymous):

No they do not know how to make proper questions

OpenStudy (anonymous):

Colin and Payton bought 79 shares of Stock A at $63 per share, 43 shares of Stock B at $119 per share, and a ten-year $7,500 bond with a 14.91% coupon for $7,000. Colin and Payton sold all stocks and bonds yesterday. Stock A was $78 per share and Stock B had a value of $131 per share. If neither stock paid a dividend, which investment has the highest rate of return? Stock A Stock B Bond Stock A and Bond

OpenStudy (anonymous):

For this question I It would be Stock A

jimthompson5910 (jim_thompson5910):

Stock A total cost = 79*63 = 4977 total revenue = 79*78 = 6162 Rate of return = (revenue - cost)/cost = (6162-4977)/4977 = 0.238095 Stock B total cost = 43*119 = 5117 total revenue = 43*131 = 5633 Rate of return = (revenue - cost)/cost = (5633-5117)/5117 = 0.100840 Bond total cost = 7000 total revenue = 7,000+0.1491*7500*10 = 18,182.5 Rate of return = (revenue - cost)/cost = (18182.5-7000)/7000 = 1.5975

jimthompson5910 (jim_thompson5910):

So it looks like the bond has the highest rate of return

OpenStudy (anonymous):

Piper invested $7,400 in a Certificate of Deposit, $4,980 in a corporate bond, $5,100 in a common stock, and $6,350 in a preferred stock, The COD has a rate of return of 3.4%; the corporate bond is 5.8%; the common stock’s rate of return is –1.9%; and the preferred stock has a rate of return of 4.6%. What is Piper’s weighted mean rate of return? 4.60% 3.90% 2.98% 3.09% @jim_thompson5910

OpenStudy (mathmate):

@jim_thompson @gabylovesu For Riley's problem, I get exactly $382316.08 as well, including all charges. Could Gaby check on two things: 1. typo's on the question,just in case 2. According to the notes, the seller is responsible for the doc.stamp tax on deed (perhaps only in Florida), so Riley should not have to pay the $1486.80.

OpenStudy (mathmate):

For Collin & Payton, \(at~maturity\) the bond will only pay 7500*1.1491=8618.25. So return = (8618.25-7000)/7000=23.11% It was not stated whether the bond was sold at maturity value. Even so, stock A (23.8%) still gets the highest return.

OpenStudy (mathmate):

Piper, For each investment, multiply the value of each investment by the rate of return to get the return amount in dollars. Calculate the total value of the investments and the total of the returns. Weighted mean rate = Total return / total investment

OpenStudy (mathmate):

@gabylovesu If you are interested in calculating Nevaeh's problem (the first one) using formulas instead of calculator, please let me know.

jimthompson5910 (jim_thompson5910):

@mathmate for the bond one, don't you have to include coupon payments as well?

OpenStudy (mathmate):

@jim_thompson5910 lol, I think it's a different interpretation between us. For "and a ten-year $7,500 bond with a 14.91% coupon for $7,000." I interpreted it as "a $7500 bond for $7000. The bond carries a 14.91% coupon".

jimthompson5910 (jim_thompson5910):

I'm not sure anymore, but I thought the bond would provide coupon payments for 10 years

OpenStudy (mathmate):

Neither am I. However, there is such a thing as a "zero coupon bond", although the question doesn't specify. In any case, the coupons would pay back the nominal value, which comes to be the same total amount. The thing is, the date of disposition (maturity) is not flexible. The question seems to give the idea the bond was sold at the same time as the shares.